# CAD-Unet: A Capsule Network-Enhanced Unet Architecture for Accurate Segmentation of COVID-19 Lung Infections from CT Images

### v1

@@ Abstract
Since the outbreak of the COVID-19 pandemic in 2019, medical imaging has emerged as a primary modality for diagnosing COVID-19 pneumonia. In clinical settings, the segmentation of lung infections from computed tomography images enables rapid and accurate quantification and diagnosis of COVID-19. Segmentation of COVID-19 infections in the lungs poses a formidable challenge, primarily due to the indistinct boundaries and limited contrast presented by ground glass opacity manifestations. Moreover, the confounding similarity between infiltrates, lung tissues, and lung walls further complicates this segmentation task. To address these challenges, this paper introduces a novel deep network architecture, called CAD-Unet, for segmenting COVID-19 lung infections. In this architecture, capsule networks are incorporated into the existing Unet framework. Capsule networks represent a novel network architecture that differs from traditional convolutional neural networks. They utilize vectors for information transfer among capsules, facilitating the extraction of intricate lesion spatial information. Additionally, we design a capsule encoder path and establish a coupling path between the unet encoder and the capsule encoder. This design maximizes the complementary advantages of both network structures while achieving efficient information fusion.
Finally, extensive experiments are conducted on four publicly available datasets, encompassing binary segmentation tasks and multi-class segmentation tasks. The experimental results demonstrate the superior segmentation performance of the proposed model. The code has been released at: https://github.com/AmanoTooko-jie/CAD-Unet.
@@ 1 Introduction
Over the years, various computerized methods have been proposed for lung segmentation in CT images, ranging from traditional image processing algorithms to advanced machine learning techniques. Deep learning, particularly convolutional neural networks (CNNs), has emerged as a dominant approach in medical image segmentation, demonstrating promising results in various applications.
However, the complex and variable nature of COVID-19 lesions within the lungs presents significant challenges for automatic segmentation using deep learning techniques. COVID-19, caused by the SARS-CoV-2 virus, manifests in the lungs through diverse and highly variable pathological findings. Firstly, the size of COVID-19-associated lesions can vary considerably, with some patients exhibiting small, focal areas of infection, while others display widespread, diffuse involvement of multiple lung lobes. This variability necessitates deep learning models to possess a high degree of flexibility and adaptability to accurately segment these patterns. Secondly, the shapes and patterns of the lesions also exhibit significant diversity. Ground-glass opacities (GGO) typically manifest as round or oval-shaped opacities, whereas consolidation (CON) may appear as irregular or patchy areas. In some cases, lesions may present with a mixed appearance, combining GGO with areas of consolidation or fibrosis. This heterogeneity in lesion morphology poses considerable challenges for deep learning models to consistently and accurately segment the infected regions. Finally, the location of lesions within the lungs can vary significantly, with some patients having lesions primarily in peripheral regions and others in more central locations. This spatial variability requires deep learning models to have a comprehensive understanding of lung anatomy and the potential locations of COVID-19 lesions.
@@
In summary, the main contributions of this paper are:
@@
  • We conduct extensive experiments on four publicly available datasets, including both binary segmentation tasks and multi-class segmentation tasks, and compare the proposed architecture against three baseline architectures (Unet [18], UNet++ [30], Att-Unet [14]) and four state-of-the-art COVID-19 segmentation architectures (binary segmentation: InfNet [6], SCOAT-Net [29], nCoVSegNet [11], and PDEAtt-Unet [3], multi-class segmentation: CopleNet [23], AnamNet [15], SCOAT-Net [29], and D-TrAttUnet [2]). The experimental results demonstrate that the proposed architecture exhibits strong segmentation performance in both binary and multi-class segmentation, outperforming the strong baselines and state-of-the-art medical image segmentation methods.
@@ 2 Related works
The COVID-19 pandemic has sparked extensive research in the field of medical imaging, particularly in the domain of lung CT lesion segmentation. There are various methods for lung CT lesion segmentation. In this section, we will only discuss segmentation architectures based on Unet and the utilization of capsule networks for target segmentation.
@@ 2.1 Segmentation architecture based on Unet
In the context of the COVID-19 pandemic, numerous researchers have adopted the Unet architecture for segmenting lesion regions in lung CT images, aiming to provide precise support for clinical diagnosis and disease monitoring. For instance, [4] initially applied Unet++ [30] to COVID-19 lung CT lesion segmentation. Unet++, an extension of the Unet framework, employs a series of nested and densely connected skip pathways between encoder and decoder subnetworks, further enhancing semantic relationships between them and achieving improved performance in lung CT lesion segmentation tasks. [29] reengineered the connection structure of Unet++ to propose SCOAT-Net, introducing a biologically motivated attention learning mechanism. This approach introduces specially designed spatial-wise and channel-wise attention modules, which collaborate to enhance the network’s attention learning and consequently improve the segmentation accuracy. Building upon Unet, [3] developed PDAtt-Unet, a pyramid dual-decoder Att-Unet [10] architecture that utilizes pyramid structures and attention gates to maintain global spatial awareness across all encoding layers. During the decoding phase, PDAtt-Unet incorporates two separate decoders employing attention gates to simultaneously segment infections and the lung region.
@@ 2.2 Capsule networks for object segmentation
Capsule networks, introduced by [19], have garnered significant attention within the field of object segmentation due to their unique architecture and potential to address certain limitations of traditional CNNs. In contrast to CNNs, that rely on pooling and hierarchical feature extraction, capsule networks aim to model the hierarchical relationships between features by representing them in the form of capsules.
@@
However, the original architecture of capsule networks and dynamic routing algorithm incurs significant computational expenses in terms of memory and runtime [8]. In order to address this concern, [8] extended the concept of convolutional capsules and redefined the dynamic routing algorithm in two crucial ways. Firstly, children capsules are selectively routed to parents capsules within a localized spatial kernel. Secondly, transformation matrices are shared among all members of the grid within a specific capsule type, yet not shared across different capsule types. Subsequently, a multitude of researchers have delved deeper into the exploration of capsule network models for the purpose of object segmentation. For instance, [20] developed the Multi-SegCaps model, EM-routing SegCaps model, and U-Net model capable of segmenting an arbitrary number of classes. These models were applied to the segmentation of individual 2D slices, utilizing two neighboring slices on each side (five in total). [21] proposed the 3DConvCaps model, which employs capsule networks for 3D image segmentation. 3DConvCaps constitutes a 3D encoder-decoder network with a convolutional capsule encoder, facilitating the learning of lower-level features using convolutional layers (short-range attention), while simultaneously modeling higher-level features (long-range dependencies) with capsule layers.
In this paper, the integration of the two techniques has been undertaken with the aim of COVID-19 lung CT lesion segmentation. However, distinct from these previous investigations, the CAD-Unet establishes a path coupling between Unet-encoder and convolutional capsules, thereby fusing the strengths of convolution and capsules to extract target features. Moreover, the incorporation of attention gates and a dual decoder has been implemented. These modules collectively interact to significantly bolster the overall performance of the CAD-Unet architecture. Subsequently, a comprehensive elucidation of the proposed network architecture is furnished.
@@ 3.2 Convolutional capsule layer
Capsule networks, originally proposed by [19], consist of multiple capsule layers and innovatively represent the information transfer between capsule layers as vectors rather than scalars. These vectors not only encode specific entity types but also describe how entities are instantiated, including their postures, textures, deformations, and the presence of these features themselves [5]. Among them, the norm of the vector indicates the probability of the entity’s existence, while the orientation of the vector indicates the configuration of the entity. Information between capsule layers is transmitted through a dynamic routing algorithm, which iteratively updates the weights between capsules, enabling higher-level capsules to effectively route the outputs of lower-level capsules. As a result, more accurate and robust feature representations are obtained. This dynamic routing mechanism is a key characteristics of capsule networks, allowing the network to learn hierarchical representations and spatial relationships of objects.
Using capsule-based networks for object segmentation poses several issues. The inherent complexity of the original capsule network architecture and dynamic routing algorithm results in significant computational demands, affecting memory usage and runtime performance [8]. To address these issues, [8] proposed SegCaps, where they introduced capsule sharing for each category, effectively reducing the computational overhead. Specifically, as illustrated in 2, the input feature map size of the capsule layer is $H\times W\times C\times A$, where $H\times W$ represents the feature map dimensions, $C$ is the number of capsule types, and $A$ is the size of each capsule. Firstly, the capsule $u_{i}$ for class $i$ is linearly mapped to a higher-level feature $\hat{u}_{i}$ using the same transformation matrix $W$. Then, all $\hat{u}_{i}$ are weighted and summed to obtain the input $s$ of the higher-level capsule. Finally, the squash function is applied to $s$ to ensure that the vector direction remains unchanged while compressing the vector length between 0 and 1, representing the probability of entity presence. The dynamic routing process can be represented (1) as follows:
@@ 3.3 Two parallel encoder pathways
Specifically, in the encoder stage of the network, we design two parallel paths: the capsule encoder path and the Unet encoder path, and the feature information from these two paths interacts with each other. The input image $x\in R^{H\times W\times C}$ is fed into both paths, where $H$ and $W$ represent the height and width of the input image, respectively, and $C$ denotes the number of channels in the input image. Now, we will provide a detailed explanation of each path.
@@
    Capsules encoder pathway: In the capsule encoding pathway, we initially generate primary capsules through a convolutional layer with a kernel size of $5\times 5$, resulting in capsules of a single category $C_{0}\in\mathbb{R}^{H\times W\times 1\times 16}$.The primary capsule $C_{0}$ is then propagated to the primary capsule layer, which is a conventional convolutional capsule layer. It undergoes one routing iteration and returns 16-dimensional capsules for two categories, denoted as $C_{1}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 2\times 16}$. Following this, at each stage, the convolutional capsule layer is coupled with the results of different levels of the Unet pathway, gradually extracting higher-level semantic information, allowing the model to effectively comprehend the complex structure of input data. Specifically, $C_{1}$ is coupled with the first layer of the Unet pathway, the coupled result is then forwarded to the secondary capsule layer, comprising three routing iterations. The output consists of 32-dimensional capsules for two categories, denoted as $C_{2}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2\times 32}$, with capsule dimensions expanding similarly through convolutional capsule layer operations. Following this, $C_{2}$ is coupled with the results of the Unet pathway and transmitted to the tertiary capsule layer, returning 32-dimensional capsules for four categories, represented as $C_{3}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4\times 32}$. Finally, $C_{3}$ is coupled with the Unet pathway results and transmitted to the quaternary capsule layer, yielding 64-dimensional capsules for four categories, denoted as $C_{4}\in\mathbb{R}^{\frac{H}{14}\times\frac{W}{14}\times 4\times 64}$. It is observable that the size of the feature map is reduced by half and the number of channels is doubled after the convolutional capsule layer. The convolutional operations within the capsule layer control the size of the output capsules, with a stride of 2 implemented for achieving a twofold downsampling. This approach enables the reduction of spatial dimensions and gradual increase in channel numbers in the feature map, facilitating the model in learning abstract hierarchical features. This process is expressed by equation (2):
    $C_{0}$ $=Conv_{5\times 5}(in),$ (2)
@@
    where $Conv_{5\times 5}(\cdot)$ is the convolution operation with a $5\times 5$ kernel, $ConvCap(\cdot)$ is the convolutional capsule layer, and $in$ is the initial input to the network.
@@
    Unet encoder pathway: In the Unet encoder pathway, we employ a Unet baseline architecture enhanced with ResBlocks [2]. This architecture comprises layers, each consisting of a ResBlock and a downsampling module. The initial layer (Layer 0) receives the original network input, while subsequent layers receive the output of the previous layer and the pathway coupling results from the capsules pathway. Specifically, we perform point-wise multiplication between the predictions from each layer of the capsules pathway and the feature extraction from each layer of the Unet pathway. This strategic fusion enables the Unet pathway to seamlessly integrate features extracted from the capsules pathway, thereby enriching its contextual understanding. Let $out_{l-1}$ represent the output of the previous layer, and the input to the current layer, denoted as $in_{l}$, is computed as $in_{l}=cat(out_{l-1},out_{l-1}\cdot C_{l})$. Here, $in_{l}$ signifies the input to the lst layer, $out_{l-1}$ is the output of the (l-1)-th layer, $C_{l}$ represents the output of the 1st capsule layer, and cat denotes the concatenation operation along the channel dimension. Following this, this input undergoes a ResBlock to extract features, followed by passage through a max-pooling layer for downsampling, resulting in the output of this layer, as depicted in equation (3):
    ${in}_{0}$ $=ResBlock(in),out_{0}=MP(in_{0}),$ (3)
    $in_{1}$ $=Cat(C_{1}\cdot out_{0},out_{0}),out_{1}=MP(RB(in_{1})),$
    $in_{2}$ $=Cat(C_{2}\cdot out_{1},out_{1}),out_{2}=MP(RB(in_{2})),$
    $in_{3}$ $=Cat(C_{3}\cdot out_{2},out_{2}),out_{3}=MP(RB(in_{3})),$
    ${in}_{4}$ $=Cat(C_{4}\cdot out_{3},out_{3}),out_{4}=RB(in_{4}),$

### v2

@@ Abstract
Since the outbreak of the COVID-19 pandemic in 2019, medical imaging has emerged as a primary modality for diagnosing COVID-19 pneumonia. In clinical settings, the segmentation of lung infections from computed tomography images enables rapid and accurate quantification and diagnosis of COVID-19. Segmentation of COVID-19 infections in the lungs poses a formidable challenge, primarily due to the indistinct boundaries and limited contrast presented by ground glass opacity manifestations. Moreover, the confounding similarity among infiltrates, lung tissues, and lung walls further complicates this segmentation task. To address these challenges, this paper introduces a novel deep network architecture, called CAD-Unet, for segmenting COVID-19 lung infections. In this architecture, capsule networks are incorporated into the existing Unet framework. Capsule networks represent a novel type of network architecture that differs from traditional convolutional neural networks. They utilize vectors for information transfer among capsules, facilitating the extraction of intricate lesion spatial information. Additionally, we design a capsule encoder path and establish a coupling path between the unet encoder and the capsule encoder. This design maximizes the complementary advantages of both network structures while achieving efficient information fusion. Finally, extensive experiments are conducted on four publicly available datasets, encompassing binary segmentation tasks and multi-class segmentation tasks. The experimental results demonstrate the superior segmentation performance of the proposed model. The code has been released at: https://github.com/AmanoTooko-jie/CAD-Unet.
@@ 1 Introduction
Over the years, various computerized methods have been proposed for lung segmentation in CT images, ranging from traditional image processing algorithms to advanced machine learning techniques. Deep learning, particularly convolutional neural networks (CNNs), has emerged as a dominant approach for medical image segmentation, demonstrating promising results in various applications.
However, the complex and variable nature of COVID-19 lesions within the lungs presents significant challenges for automatic segmentation using deep learning techniques. COVID-19, caused by the SARS-CoV-2 virus, manifests in the lungs through diverse and highly variable pathological findings. Firstly, the size of COVID-19-associated lesions can vary considerably, with some patients exhibiting small, focal areas of infection, while others display widespread, diffuse involvement of multiple lung lobes. This variability necessitates deep learning models to possess a high degree of flexibility and adaptability to accurately segment these patterns. Secondly, the shapes and patterns of the lesions also exhibit significant diversity. GGOs typically manifest as round or oval-shaped opacities, whereas CONs may appear as irregular or patchy areas. In some cases, lesions may present with a mixed appearance, combining GGOs with areas of consolidation or fibrosis. This heterogeneity in lesion morphology poses considerable challenges for deep learning models to consistently and accurately segment the infected regions. Finally, the location of lesions within the lungs can vary significantly, with some patients having lesions primarily in peripheral regions and others in more central locations. This spatial variability requires deep learning models to have a comprehensive understanding of lung anatomy and the potential locations of COVID-19 lesions.
@@
In summary, the main contributions of this paper are given as follows:
@@
  • We conduct extensive experiments on four publicly available datasets, including both binary segmentation tasks and multi-class segmentation tasks, and compare the proposed architecture against three baseline architectures (Unet [20], UNet++ [32], Att-Unet [16]) and four state-of-the-art COVID-19 segmentation architectures (binary segmentation: InfNet [6], SCOAT-Net [31], nCoVSegNet [13], PDEAtt-Unet [3] and D-TrAttUnet [2], multi-class segmentation: CopleNet [25], AnamNet [17], SCOAT-Net [31], and D-TrAttUnet). In addition, we also integrate our architecture into nnUNet [7] and compare it with the native nnUNet architecture. The experimental results demonstrate that the proposed architecture exhibits strong segmentation performance in both binary and multi-class segmentation, outperforming the strong baselines and state-of-the-art medical image segmentation methods.
@@ 2 Related works
The COVID-19 pandemic has sparked extensive research in the field of medical imaging, particularly in the domain of lung CT lesion segmentation. There are various methods for lung CT lesion segmentation. In this section, we will only discuss segmentation architectures based on Unet and the utilization of capsule networks for object segmentation.
@@ 2.1 Segmentation architecture based on Unet
In the context of the COVID-19 pandemic, numerous researchers have adopted the Unet architecture for segmenting lesion regions in lung CT images, aiming to provide precise support for clinical diagnosis and disease monitoring. For instance, [4] initially applied Unet++ [32] to COVID-19 lung CT lesion segmentation. Unet++, an extension of the Unet framework, employs a series of nested and densely connected skip pathways between the encoder and decoder subnetworks, further enhancing the semantic relationships between them and achieving improved performance in lung CT lesion segmentation tasks. [31] reengineered the connection structure of Unet++ to propose SCOAT-Net, introducing a biologically motivated attention learning mechanism. This approach introduces specially designed spatial-wise and channel-wise attention modules, which collaborate to enhance the network’s attention learning and consequently improve the segmentation accuracy. Building upon Unet, [3] developed PDAtt-Unet, a pyramid dual-decoder Att-Unet [12] architecture that utilizes pyramid structures and attention gates to maintain global spatial awareness across all encoding layers. During the decoding phase, PDAtt-Unet incorporates two separate decoders employing attention gates to simultaneously segment infections and the lung region.
@@ 2.2 Capsule networks for object segmentation
Capsule networks, introduced by [21], have attracted considerable attention in the domain of object segmentation, owing to their distinctive architecture and capacity to mitigate specific constraints associated with conventional CNNs. Unlike CNNs, which depend on pooling and hierarchical feature extraction, capsule networks endeavor to characterize the hierarchical interdependencies among features by encapsulating them within capsules.
@@
However, the original architecture of capsule networks and the dynamic routing algorithm incur significant computational expenses in terms of memory and runtime [10]. To address this concern, [10] extended the concept of convolutional capsules and redefined the dynamic routing algorithm in two crucial ways. Firstly, children capsules are selectively routed to parents capsules within a localized spatial kernel. Secondly, transformation matrices are shared among all members of the grid within a specific capsule type, but not shared across different capsule types. Subsequently, a multitude of researchers have delved deeper into the exploration of capsule network models for the purpose of object segmentation. For instance, [22] developed the Multi-SegCaps model, EM-routing SegCaps model, and U-Net model capable of segmenting an arbitrary number of classes. These models were applied to the segmentation of individual 2D slices, utilizing two neighboring slices on each side (five in total). [23] proposed the 3DConvCaps model, which employs capsule networks for 3D image segmentation. 3DConvCaps constitutes a 3D encoder-decoder network with a convolutional capsule encoder, facilitating the learning of lower-level features using convolutional layers (short-range attention), while simultaneously modeling higher-level features (long-range dependencies) with capsule layers.
In this paper, the integration of the two techniques has been undertaken with the aim of COVID-19 lung CT lesion segmentation. However, distinct from these previous investigations, the CAD-Unet establishes a path coupling between the Unet-encoder and convolutional capsules, thereby fusing the strengths of convolution and capsules to extract target features. Moreover, the incorporation of attention gates and a dual decoders has been implemented. These modules collectively interact to significantly bolster the overall performance of the CAD-Unet architecture. Subsequently, a comprehensive elucidation of the proposed network architecture is furnished.
@@ 3.2 Convolutional capsule layer
Capsule networks consist of multiple capsule layers and innovatively represent the information transfer between capsule layers as vectors rather than scalars. These vectors not only encode specific entity types but also describe how entities are instantiated, including their postures, textures, deformations, and the presence of these features themselves [5]. Among them, the norm of the vector indicates the probability of the entity’s existence, while the orientation of the vector indicates the configuration of the entity. Information between capsule layers is transmitted through a dynamic routing algorithm, which iteratively updates the weights between capsules, enabling higher-level capsules to effectively route the outputs of lower-level capsules. As a result, more accurate and robust feature representations are obtained. This dynamic routing mechanism is a key characteristics of capsule networks, allowing the network to learn hierarchical representations and spatial relationships of objects.
Using capsule-based networks for object segmentation poses several issues. The inherent complexity of the original capsule network architecture and dynamic routing algorithm results in significant computational demands, affecting memory usage and runtime performance [10]. To address these issues, [10] proposed SegCaps, which introduces capsule sharing for each category, effectively reducing the computational overhead. Specifically, as illustrated in Fig.2, the input feature map size of the capsule layer is $H\times W\times C\times A$, where $H\times W$ represents the feature map dimensions, $C$ is the number of capsule types, and $A$ is the size of each capsule. Firstly, the capsule $u_{i}$ for class $i$ is linearly mapped to a higher-level feature $\hat{u}_{i}$ using the same transformation matrix $W$. Then, all $\hat{u}_{i}$ are weighted and summed to obtain the input $s$ of the higher-level capsule. Finally, the squash function is applied to $s$ to ensure that the vector direction remains unchanged while compressing the vector length between 0 and 1, representing the probability of entity presence. The dynamic routing process can be represented (1) as follows:
@@ 3.3 Two parallel encoder pathways
Specifically, in the encoder stage of the network, we design two parallel paths: the capsule encoder path and the Unet encoder path, and the feature information from these two paths interacts with each other. The input image $x\in\mathbb{R}^{H\times W\times C}$ is fed into both paths, where $H$ and $W$ represent the height and width of the input image, respectively, and $C$ denotes the number of channels in the input image. Now, we will provide a detailed explanation of each path.
@@
    Capsules encoder pathway: In the capsule encoding pathway, we initially generate primary capsules through a convolutional layer with a kernel size of $5\times 5$, resulting in capsules of a single category $C_{0}\in\mathbb{R}^{H\times W\times 1\times 16}$.The primary capsule $C_{0}$ is then propagated to the primary capsule layer, which is a conventional convolutional capsule layer. It undergoes one routing iteration and returns 16-dimensional capsules for two categories, denoted as $C_{1}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 2\times 16}$. Following this, at each stage, the convolutional capsule layer is coupled with the results of different levels of the Unet pathway, gradually extracting higher-level semantic information, allowing the model to effectively comprehend the complex structure of input data. Specifically, $C_{1}$ is coupled with the first layer of the Unet pathway, the coupled result is then forwarded to the secondary capsule layer, comprising three routing iterations. The output consists of 32-dimensional capsules for two categories, denoted as $C_{2}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2\times 32}$, with capsule dimensions expanding similarly through convolutional capsule layer operations. Following this, $C_{2}$ is coupled with the results of the Unet pathway and transmitted to the tertiary capsule layer, returning 32-dimensional capsules for four categories, represented as $C_{3}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4\times 32}$. Finally, $C_{3}$ is coupled with the Unet pathway results and transmitted to the quaternary capsule layer, yielding 64-dimensional capsules for four categories, denoted as $C_{4}\in\mathbb{R}^{\frac{H}{14}\times\frac{W}{14}\times 4\times 64}$. It is evident that the size of the feature map is halved, while the number of channels is doubled after the convolutional capsule layer. The convolutional operations within the capsule layer regulate the size of the output capsules, with a stride of 2 employed to achieve twofold downsampling. This strategy reduces the spatial dimensions and incrementally increases the channel numbers in the feature map, thereby enabling the model to learn abstract hierarchical features. This process is formalized in equation (2):
    $C_{0}$ $=Conv_{5\times 5}(x_{in}),$ (2)
@@
    where $Conv_{5\times 5}(\cdot)$ is the convolution operation with a $5\times 5$ kernel, $ConvCap(\cdot)$ is the convolutional capsule layer, and $x_{in}$ is the initial input to the network.
@@
    Unet encoder pathway: In the Unet encoder pathway, we employ a Unet baseline architecture enhanced with ResBlocks [2]. This architecture comprises multiple layers, each consisting of a ResBlock and a downsampling module. The initial layer (Layer 0) receives the original network input, while subsequent layers receive the output of the previous layer and the pathway coupling results from the capsules pathway. Specifically, we perform point-wise multiplication between the predictions from each layer of the capsules pathway and the feature extraction from each layer of the Unet pathway. This strategic fusion enables the Unet pathway to seamlessly integrate features extracted from the capsules pathway, thereby enriching its contextual understanding. Let $out_{l-1}$ represent the output of the previous layer, and the input to the current layer, denoted as $in_{l}$, is computed as $in_{l}=cat(out_{l-1},out_{l-1}\cdot C_{l})$. Here, $in_{l}$ signifies the input to the $l$-th layer, $out_{l-1}$ is the output of the ($l-1$)-th layer, $C_{l}$ represents the output of the $l$-th capsule layer, and $Cat(\cdot)$ denotes the concatenation operation along the channel dimension. Following this, this input undergoes a ResBlock to extract features, followed by passage through a max-pooling layer for downsampling, resulting in the output of this layer, as depicted in equation (3):
    ${in}_{0}$ $=ResBlock(in),~{}out_{0}=MP(in_{0}),$ (3)
    $in_{1}$ $=Cat(C_{1}\cdot out_{0},out_{0}),~{}out_{1}=MP(RB(in_{1})),$
    $in_{2}$ $=Cat(C_{2}\cdot out_{1},out_{1}),~{}out_{2}=MP(RB(in_{2})),$
    $in_{3}$ $=Cat(C_{3}\cdot out_{2},out_{2}),~{}out_{3}=MP(RB(in_{3})),$
    ${in}_{4}$ $=Cat(C_{4}\cdot out_{3},out_{3}),~{}out_{4}=RB(in_{4}),$